Elephant2020:

state the number of terms in each of the following expressions : i) 2a-b ii) 3 x x + a/2 iii) 3x - x/p iv) a divided by x x b + c v) 3 x divided by 2 + y + 4 can u just post this as a question? @rainbowsalt

rainbowsalt:

(i) 2a- b = 2 (ii) 3 x x + \[\frac{ a }{ 2 }\] = 3x + \[\frac{ a}{ 2}\] = 2 (iii) 3x - \[\frac{ x}{ p }\] = 2 (iv) a divided x x b+c = \[\frac{ ab }{ x }\] + c = 2 (v) 3x divided 2+y + 4 = \[\frac{ 3x }{ 2 }\] + y+ 4 = 3

Lifesuck:

Great job

rainbowsalt:

ty

Lifesuck:

Ur yw

BrightStar:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @rainbowsalt (i) 2a- b = 2 (ii) 3 x x + \[\frac{ a }{ 2 }\] = 3x + \[\frac{ a}{ 2}\] = 2 (iii) 3x - \[\frac{ x}{ p }\] = 2 (iv) a divided x x b+c = \[\frac{ ab }{ x }\] + c = 2 (v) 3x divided 2+y + 4 = \[\frac{ 3x }{ 2 }\] + y+ 4 = 3 \(\color{#0cbb34}{\text{End of Quote}}\) is this copied from another web...?

rainbowsalt:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @BrightStar \(\color{#0cbb34}{\text{Originally Posted by}}\) @rainbowsalt (i) 2a- b = 2 (ii) 3 x x + \[\frac{ a }{ 2 }\] = 3x + \[\frac{ a}{ 2}\] = 2 (iii) 3x - \[\frac{ x}{ p }\] = 2 (iv) a divided x x b+c = \[\frac{ ab }{ x }\] + c = 2 (v) 3x divided 2+y + 4 = \[\frac{ 3x }{ 2 }\] + y+ 4 = 3 \(\color{#0cbb34}{\text{End of Quote}}\) is this copied from another web...? \(\color{#0cbb34}{\text{End of Quote}}\) nope u can search google

BrightStar:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @rainbowsalt \(\color{#0cbb34}{\text{Originally Posted by}}\) @BrightStar \(\color{#0cbb34}{\text{Originally Posted by}}\) @rainbowsalt (i) 2a- b = 2 (ii) 3 x x + \[\frac{ a }{ 2 }\] = 3x + \[\frac{ a}{ 2}\] = 2 (iii) 3x - \[\frac{ x}{ p }\] = 2 (iv) a divided x x b+c = \[\frac{ ab }{ x }\] + c = 2 (v) 3x divided 2+y + 4 = \[\frac{ 3x }{ 2 }\] + y+ 4 = 3 \(\color{#0cbb34}{\text{End of Quote}}\) is this copied from another web...? \(\color{#0cbb34}{\text{End of Quote}}\) nope u can search google \(\color{#0cbb34}{\text{End of Quote}}\) kk